### v1

@@ 4 Conclusion
We have shown that tile assembly models incorporating a pre-determined sequence linked by a backbone can deterministically assemble all achievable shapes – those with a Hamilton path – using a finite universal assembly kit of tiles, unlike the conventional temperature-1 aTAM. This fact holds true even if we do not permit neutral interactions between faces. We have also presented evidence that the backboned aTAM needs fewer tile types than the conventional temperature-1 aTAM to deterministically assemble a single target shape.
@@
We have also explored the roles of both the pre-specified sequence and the geometrical constraints imposed by the backbone in guiding assembly. Through our results on the sequenced aTAM, we showed that keeping the sequence encoding of tiles but removing geometric restrictions on the positions of added tiles allows a large space of possible shapes, but excludes the possibility of a universal assembly kit for these shapes. Our results tentatively suggest that the backbone constraint not only restricts the space of shapes that can be assembled, but also actively facilitates deterministic assembly with a small tile set, since ambiguities of where to place tiles can be avoided. A key question that remains open is whether the sequenced aTAM can deterministically assemble any shape accessible to the backboned aTAM using a finite assembly kit, and if so, how the minimal complexity of this finite assembly kit compares to that of the backboned aTAM.
From an informational perspective, it is interesting to look at the efficiency with which shapes are specified by sequences. To encode all shapes within a space, it is necessary that the combinatorial possibilities of the sequence grow a least as fast with sequence length $N$ as the number of accessible shapes. For the backboned aTAM with neutral interactions, we have identified a universal assembly kit that almost reaches this limit; the shape space scales as $O(2.638^{N})$, implying at least 3 distinct tiles are required, and the universal assembly kit contains only 4 tiles, one of which is only used at the start of a sequence. For the model excluding neutral interactions, the universal assembly kit we have identified is much less efficient, as it contains 208 tiles. With this kit, many sequences must form identical structures, others will be non-deterministic and some will be prematurely terminated. These contributions will be even more dominant for the sequenced aTAM, whose shape space scales as $O(4.5252^{N})$ but for which no universal assembly kit exists. The extent to which the different factors contribute to inefficiency in specifying a shape with a sequence are an intriguing unresolved question.
@@
This highly abstract study was initially motivated by the biologically relevant question of what a pre-formed backbone contributes to the self-assembly of RNA and proteins. Our results suggest that, at least in simplified models, the presence of a backbone has a qualitative, rather than merely quantitative, effect on the number of structures that can be reliably formed with a finite set of building blocks, and that both the sequence and the geometric constraints imposed by the backbone are important in directing successful assembly. These results are likely to generalise beyond the specific details of the model considered here, but it would also be instructive to explore more realistic models. In particular, biological assembly occurs at finite temperature with interactions of moderate strength. Contacts can form and break, and cooperativity between units is essential in forming long-lasting bonds. Indeed, the need for cooperative interactions may explain how RNA folding can operate with only 4 unique bases, and protein folding can occur with only 20 different side chains; in practice, cooperatively interacting domains likely provide a much larger effective set of building blocks to direct assembly. Testing this hypothesis with a model of backboned assembly that incorporates finite interaction strengths and cooperative bonding would be a natural next step.
    Acknowledgments We thank Ard Louis, Jordan Juritz and Benjamin Qureshi for their helpful discussions.
    Declarations
@@
    Theorem A.28 below replaces Theorem 3.11 in the main text for the case where repulsive interactions are permitted. Theorem A.28 has additional assumptions on the properties of the shape in question relative to Theorem 3.11, but the treeangle shape defined in the main text was constructed to satisfy this extended set of assumptions. Hence, Theorem A.28 leads directly into Lemma 3.15 in the main text.
    Now, we extend the arguments from Theorem 3.11 for cases where the strength function $g$ is allowed to take on the values $g<0$. The only difference is we now need to consider tile placement being blocked by repulsion, rather than just overlap of tiles, which could in principle result in the growth of distinct shapes from the same glue type. To understand how this may happen, consider the following set up. Let $\Psi$ be a complete trajectory of $\mathcal{P}$, with entries $A_{t}=A_{0,t}+A_{1,t}+A_{2,t}+...+A_{M,t}$ where $E(A_{i})=S_{i}$ and $A_{i,t}$ is the subconfiguration of $A_{i}$ appearing at time $t$ in $\Psi$. Assuming that two ‘growth’ faces have the same glue $A_{0}(\ddot{z}_{i},\ddot{k}_{i})=A_{0}(\ddot{z}_{j},\ddot{k}_{j})$, using the approach in Lemma 3.9, we can construct another ‘partial’ trajectory $\Psi^{\prime}$ with the subconfigurations $A_{i,t}$ and $A_{j,t}$ swapped that stops at the first time point $t_{s}$ when repulsion due to an interaction with $g<0$ precludes a tile from being added to an empty coordinate for the first time. As with Theorem 3.9, we can write $A_{t}^{\prime}=A_{0}+A_{i\rightarrow j,t}+A_{j\rightarrow i,t}+A_{c,t}$ as the entries of the trajectory $\Psi^{\prime}$. (As a reminder, $A_{j\rightarrow i}$ represents the configuration $A_{j}$ transformed so that it now ‘grows’ from $(\ddot{z}_{i},\ddot{k}_{i})$ instead).
    For this trajectory $\Psi^{\prime}$, at some time $t_{s}$, a single tile configuration $a_{i}^{des}$ with tile consistent with entry $t_{s}$ of $Q$ forms an attractive interaction with configuration $A_{j\rightarrow i,t_{s}}$, but is prevented from being added because it forms some repulsive interaction with another tile. We call $a_{i}^{des}$ a destabilizing tile, and its corresponding coordinate $z_{i}^{des}$ a destabilizing coordinate. The trajectory $\Psi^{\prime}$ can then proceed, and eventually terminate either prematurely or when the last tile of $Q$ is reached, thus $\Psi^{\prime}$ is a complete trajectory. In this complete trajectory, say $A_{j\rightarrow i,t_{s}}$ grows into a configuration $A_{j\rightarrow i}^{\prime}$. As a result of the tile blocking event arising from $a_{i}^{des}$, it is possible that deterministic growth can occur that results in $E(A_{j\rightarrow i}^{\prime})=E(A_{i})\not\simeq E(A_{j})$ and $E(A_{i\rightarrow j}^{\prime})=E(A_{j})\not\simeq E(A_{i})$. Repulsion-derived tile blocking can thus cause the reasoning in Theorem 3.11 to fail.
@@
      For any $i\in{1,..,M}$, $G_{S}(S_{i})$ is a tree with at least one branching node, and with branching points in $S_{i}$ labelled $v_{i,n}$. Then, every subtree subshape with origin at some $v_{i,n}$ has a unique shape (Figure 16).
@@
    Having expanded our assumptions on our target shape, we wish to consider when repulsion-induced blocking may or may not be significant for our arguments. Consider a sequenced aTAM instance $\mathcal{P}=(A_{0},Q,g)$ that produces a target shape $S$ from starting shape that $S_{0}=E(A_{0})$. The terminal configuration $A_{\omega}$ of a trajectory of $\mathcal{P}$ has subconfigurations $A_{i}$ such that $E(A_{i})=S_{i}$. Assume that partial sub-configurations $A_{j,t_{s}}$ and $A_{i,t_{s}}=A_{j\rightarrow i,t_{s}}$ could be reached during complete trajectories at some time $t_{s}$. We say that $j$ and $i$ are differentially blocked if, for some single tile configuration $a_{j}$, $A_{j,t_{s}}+a_{j}$ can be reached during a complete trajectory but $A_{i,t_{s}}+a_{j\rightarrow i}$ cannot (or vice versa), due to repulsion-induced blocking (Hence $a_{i}^{des}=a_{j\rightarrow i}$ is the destabilizing tile with coordinate $z_{i}^{des}$). In the following lemma, we show that differential blocking is essential to break the arguments in Theorem 3.11.
    Lemma A.22.
@@
    Proof A.23.
    Consider whether it is possible to detemrinistically produce an assembly with $S_{i}\not\simeq S_{j}$ if $S_{i}$ and $S_{j}$ are anchored to the same glue type. First, if the addition of a tile to partial subconfigurations corresponding to $i$ and $j$ is never blocked due to some repulsive interaction $g<0$, then Theorem 3.11 is sufficient to argue that $S_{i}\simeq S_{j}$ necessarily, since the assumptions of Definition 7 are a subset of those of Theorem A.28 and the contradictions that arise in Theorem A.28 remain upheld.
    We now allow for non-differential repulsion-induced tile blocking. Consider (without loss of generality) a non-differential repulsion-induced tile block that occurs when trying to add a tile to a partial configuration $A_{i,t_{s}}$, where partial subconfiguration $A_{j,t_{s}}=A_{i\rightarrow j,t_{s}}$ can also be reached in a complete trajectory. This non-differential tile block does not result in partial subconfigurations $A_{j,t_{s}^{\prime}}$ whose equivalents $A_{i,t_{s}^{\prime}}=A_{j\rightarrow i,t_{s}^{\prime}}$ cannot be reached in a complete trajectory. Unless differential tile blocking occurs elsewhere, the assumed properties of the target shape then imply that for all partial subconfigurations of $A_{j}$ that can be reached in a complete trajectory, the equivalent partial subconfiguration of $A_{i}$ can also be reached and the proof of Theorem A.28 still applies. Either $S_{i}\simeq S_{j}$ and a violation of assumption 3 of our target shape arises, or growth must be non-deterministic.
    Hence, in the absence of differential blocking, each $S_{i}$ must be anchored to a unique glue type, so at least $M$ glue types are required in $A_{0}$. Since each tile type can contain at most $4$ unique glue types, at least $\frac{M}{4}$ tile types, are needed in $A_{0}$ to assemble $S$ deterministically.
    Lemma A.22 has the following corollary, which states that we would still need $M/4$ tile types if tile blocking only occurs within given subshapes $S_{i}$ rather than in between them.
    Corollary A.24.
@@
    Proof A.25.
    We once again assume that $S_{i}$ and $S_{j}$ are anchored to the same glue type. Consider (without loss of generality) a tile blocking event in which $z_{i}^{des}$ only neighbours $S_{i}$, and that the tile is being added to the partial subconfiguration $A_{i,t_{s}}$. Unless a previous differential blocking event has occurred, from the properties of the target shape, the equivalent partial subconfiguration $A_{j,t_{s}}=A_{i\rightarrow j,t_{s}}$ can also be reached in a complete trajectory if assembly is to be oriented-deterministic. $z_{i\rightarrow j}^{des}$ would therefore also experience repulsion-induced blocking, and hence blocking events in which the destabilizing coordinate $z_{i}^{des}$ only neighbours one $S_{i}$ are non-differential. This Corollary then follows directly as a result of Lemma A.22.
    Lemma A.26.
@@
    First, assume such a tile blocking event occurs, but $z_{i}^{des}$ is subsequently filled in at some later time (in the trajectory Ψsuperscriptnormal-$\Psi^{\prime}$). Then $S_{i}$ would neighbour $S_{k}$, and hence a contradiction arises with either assumption 1 or assumption 2. Hence, we only need to consider differential tile blocks where here $z_{i}^{des}$ is adjacent to $S_{i}$ as well as some $S_{k}$ with $i\neq k$, and $z_{i}^{des}$ remains unoccupied in $A_{\omega}^{\prime}$, the terminal configuration of Ψsuperscriptnormal-$\Psi^{\prime}$.
    We now proceed to argue that in this final case, features of $A_{\omega}^{\prime}$ result in a final shape $S$ that necessarily breaks one of the assumptions 2-6 of Definition 7. $a^{des}_{i}$ neighbours some tile $a^{att}_{j\rightarrow i}$ in $A_{j\rightarrow i,t_{s}}$, onto which it may be attracted, and some other tile $a^{rpl}$ that blocks its addition. Hence, $a^{att}_{j\rightarrow i}$ and $a^{rpl}$ must be of distance $2$ from each other. The assumptions on $S$ that we have laid out forbid points belonging to any two distinct $S_{i}$ and $S_{j}$ from being within distance 2 of each other, with exceptions around the vicinity of junctions where $S_{0}$ meets some $S_{i}$. More specifically, in order to fulfill assumptions 4 and 5, $a^{att}_{j\rightarrow i}$ can only occupy the position $\dot{z}_{i}$ or some neighbour of $\dot{z}_{i}$.
@@
    Hence, a contradiction in one of our assumptions arises if a differential block with $z_{i}^{des}$ that neighbours $S_{i}$ and $S_{k}$ with $k\neq i$ occurs.
    Theorem A.28.
@@
    Assumption 7 and determinism imply that $\mathcal{P}$ is oriented-deterministic. As per Lemma A.22 and Corollary A.24, only differential blocking where $z_{i}^{des}$ neighbours $S_{i}$ and $S_{k}$ with $k\neq i$ will allow $A_{0}$ to require fewer than $\frac{M}{4}$ tile types. However, Lemma A.26 shows that such a differential tile block necessarily results in $S$ failing one of assumptions 2- 6 in Definition 7. Hence, a target shape fulfilling the assumptions of Definition 7 cannot be deterministically assembled by $\mathcal{P}$ if $A_{0}$ has fewer than $\frac{M}{4}$ tile types.
@@
    The treeangle shape defined in the main text was constructed such that it can be partitioned in a way that satisfies the assumptions of Definition 7. Hence this Theorem can be used in the proof of Lemma 3.15, taking the place of Theorem 3.11 which is only valid if $g\geq 0$.
@@
    Figure 16: An illustration of assumption 3 in the definition of shapes $S_{i}$ used in Theorem A.28. The two shapes are two instances of $S_{i}$, where the lower white tiles connect to some larger $S_{0}$. A single branching point (also the origin tile of the subtree subshapes) is given in orange, while root tiles are outlined in red. The left oriented shape obeys assumption 3, as each subtree subshape (coloured in blue, yellow and green) are distinct, while the right oriented shape violates this assumption as the yellow rooted subtree subshape is equivalent to the green under a $90^{\circ}$ clockwise rotation followed by a $(1,-1)$ translation.
@@
    Figure 17: Illustrations of assumptions 4 and 5 in the definition of shapes in Theorem A.28. Tiles belonging to $S_{0}$ are in white, and tiles belonging to each $S_{i}$ are assigned a single color. Tiles violating assumptions are outlined in red. a. An illustration of how assumption 4 can be violated. Violation 1 is due to a tile of $S_{j}$ (with $\ddot{z}_{j}=\ddot{z}_{i}$) that is not $\dot{z}_{j}$ being within distance $2$ of $\dot{z}_{i}$, while violation 2 is due to a tile of $S_{k}$ ($\ddot{z}_{k}\neq\ddot{z}_{i}$) being within distance $2$ of an arbitrary tile of $S_{i}$. b. An illustration of how assumption 5 can be violated. Violation 1 is due to a tile of $S_{0}$ being within distance 2 of a tile of $S_{i}$, with both tiles being far from $\ddot{z}_{i}$ and $\dot{z}_{i}$. Violation 2 is due to two tiles (the violating tile and the purple outlined tile) in $S_{0}$ being within distance 2 of $\dot{z}_{i}$. Finally, violation $3$ is due to a tile of $S_{0}$, far from $\ddot{z}_{i}$, being within distance 2 of a neighbour of $\dot{z}_{i}$.
@@
    Figure 18: Theorem A.28 lays out two possibilities in the case where a tile-blocked coordinate $z^{des}_{i}$ remains unfilled at the end of a complete trajectory. We show here fragments of the terminal configuration $A_{\omega,c}^{\prime}$. $\mathbf{A_{j\rightarrow i}}$ is given in yellow, $\mathbf{A_{j}}$ is given in blue and destabilizing tiles $\mathbf{z_{i}^{des}}$ are in dotted red,. We illustrate how the two possibilities violate assumptions on $S$ laid out at the beginning of Theorem A.28. a. An illustration of how case 1 violates assumptions on $S$. The striped tile can be either a tile of $\mathbf{A_{0}}$ or $\mathbf{\dot{z}_{k}}$. Arrows represent some arbitrary tree-shaped configurations, where arrows of the same color represent configurations with equivalent (up to rotations and translations) shapes. The tile at $z_{j}^{nb}$ cannot be destabilized upon transformation into $z_{i}^{nb}$ as it is too far away from any coordinate of $S_{0}$ or any other $S_{k}$, and hence $z_{i}^{nb}\neq z_{i}^{des}$. Since branches that grow from $z_{i}^{nb}$ are too far away from any $S_{k}$ for $k\neq i$ to be blocked, The branches that grow from $z_{i}^{nb}$ and $z_{j}^{nb}$ must have identical shapes, violating assumptions on the shape $S$. b. An illustration of how case 2 violates assumptions on $S$. Green tiles are $\dot{z}_{k}$ for some $k\neq i,j$. The green arrows represent arbitrary tree-shaped configurations. If $\ddot{z}_{i}$ is branching (Top left), or if $z_{i}^{des}$ neighbours $\dot{z}_{i}$ (Top right), $\{\ddot{z}_{i},\dot{z}_{i}\}$ is a straight line segment of length 2, breaking assumption 6. Otherwise, if the destabilizing tile neighbours $\ddot{z}_{i}$ and a neighbour of $\dot{z}_{i}$ (Bottom left), but its equivalent is present in $S_{j}$, then $S_{j}$ has two coordinates neighbouring $S_{0}$ (Bottom right).

### v2

@@ 4 Conclusion
That sequence-directed assembly can help reduce the number of tiles required to assemble shapes is intuitively understood. Here, we show that the details of a sequence-directed assembly map have important implications on the ability of the given assembly map to assemble all shapes in its shape space with a finite set of building blocks. Through our results on the sequenced aTAM, we showed that keeping the sequence encoding of tiles but removing geometric restrictions on the positions of added tiles allows a larger space of possible shapes, but excludes the possibility of a universal assembly kit for these shapes. Our results tentatively suggest that the backbone constraint not only restricts the space of shapes that can be assembled, but also actively facilitates deterministic assembly with a small tile set, since ambiguities of where to place tiles can be avoided. So far, our results have explored rotatable but not flippable tiles, and it would be interesting to consider whether allowing tiles to flip would have any significant impacts on our results. Alternatively, taking the model into 3 dimensions would, presumably, inhibit the effectiveness of the shape-filling approach for the sequenced aTAM and may result in less contrived structures than treeangles that cannot be assembled with a finite assembly kit. Another key question that remains open is whether the sequenced aTAM can deterministically assemble any shape accessible to the backboned aTAM using a finite assembly kit (under the assumption of attractive inter-tile interactions), and if so, how the minimal complexity of this finite assembly kit compares to that of the backboned aTAM.
An initially surprising result was the degree of inefficiency with which the information in the sequence program is used to direct assembly in the models with only attractive inter-tile interactions. The mechanistic nature of the assembly process has a large effect, and the failure of many sequences to form deterministic, well-defined structures necessitates more tile types, which in turn results in redundancy, further increasing inefficiency.
This highly abstract study was initially motivated by the biologically relevant question of what a pre-formed backbone contributes to the self-assembly of RNA and proteins. Our results suggest that, at least in simplified models, the presence of a backbone has a qualitative, rather than merely quantitative, effect on the number of structures that can be reliably formed with a finite set of building blocks, and that both the sequence and the geometric constraints imposed by the backbone are important in directing successful assembly. These results are likely to generalise beyond the specific details of the model considered here. Moreover, it is interesting to note that both proteins (?) and RNA (?) have large neutral spaces, within which many sequences fold into the same structure, and also that many sequences do not produce well-defined folds (?). Both imply some degree of inefficiency in terms of converting the sequence program into structure formation.
@@
Nonetheless, it would also be instructive to explore more realistic models. In particular, biological assembly occurs at finite temperature with interactions of moderate strength. Contacts can form and break, and cooperativity between units is essential in forming long-lasting bonds. Indeed, the need for cooperative interactions may explain how RNA folding can operate with only 4 unique bases, and protein folding can occur with only 20 different side chains; in practice, cooperatively interacting domains likely provide a much larger effective set of building blocks to direct assembly. Testing this hypothesis with a model of backboned assembly that incorporates finite interaction strengths and cooperative bonding would be a natural next step.
Ethical Approval
Not applicable, as no human or animal subjects were involved in this work.
Competing Interests
@@
Funding
JG was supported by an Imperial College President’s PhD Scholarship, and TEO by a Royal Society University Research Fellowship.
Authors’ Contributions
JG and TEO planned the research. JG performed the research. JG and TEO wrote the manuscript.
Data Access Statement
Data and code may be found at 10.5281/zenodo.15020244.
We thank Ard Louis, Jordan Juritz and Benjamin Qureshi for their helpful discussions.
@@
Theorem 6 below replaces Theorem 3 in the main text for the case where repulsive interactions are permitted. Theorem 6 has additional assumptions on the properties of the shape in question relative to Theorem 3, but the treeangle shape defined in the main text was constructed to satisfy this extended set of assumptions. Hence, Theorem 6 leads directly into Lemma 3 in the main text.
Now, we extend the arguments from Theorem 3 for cases where the strength function $g$ is allowed to take on the values $g<0$. The only difference is we now need to consider tile placement being blocked by repulsion, rather than just overlap of tiles, which could in principle result in the growth of distinct shapes from the same glue type. To understand how this may happen, consider the following set up. Let $\Psi$ be a complete trajectory of $\mathcal{P}$, with entries $A_{t}=A_{0,t}+A_{1,t}+A_{2,t}+...+A_{M,t}$ where $E(A_{i})=S_{i}$ and $A_{i,t}$ is the subconfiguration of $A_{i}$ appearing at time $t$ in $\Psi$. Assuming that two ‘growth’ faces have the same glue $A_{0}(\ddot{z}_{i},\ddot{k}_{i})=A_{0}(\ddot{z}_{j},\ddot{k}_{j})$, using the approach in Lemma 2, we can construct another ‘partial’ trajectory $\Psi^{\prime}$ with the subconfigurations $A_{i,t}$ and $A_{j,t}$ swapped that stops at the first time point $t_{s}$ when repulsion due to an interaction with $g<0$ precludes a tile from being added to an empty coordinate for the first time. As with Theorem 2, we can write $A_{t}^{\prime}=A_{0}+A_{i\rightarrow j,t}+A_{j\rightarrow i,t}+A_{c,t}$ as the entries of the trajectory $\Psi^{\prime}$. (As a reminder, $A_{j\rightarrow i}$ represents the configuration $A_{j}$ transformed so that it now ‘grows’ from $(\ddot{z}_{i},\ddot{k}_{i})$ instead).
For this trajectory $\Psi^{\prime}$, at some time $t_{s}$, a single tile configuration $a_{i}^{des}$ with tile consistent with entry $t_{s}$ of $Q$ forms an attractive interaction with configuration $A_{j\rightarrow i,t_{s}}$, but is prevented from being added because it forms some repulsive interaction with another tile. We call $a_{i}^{des}$ a destabilizing tile, and its corresponding coordinate $z_{i}^{des}$ a destabilizing coordinate. The trajectory $\Psi^{\prime}$ can then proceed, and eventually terminate either prematurely or when the last tile of $Q$ is reached, thus $\Psi^{\prime}$ is a complete trajectory. In this complete trajectory, say $A_{j\rightarrow i,t_{s}}$ grows into a configuration $A_{j\rightarrow i}^{\prime}$. As a result of the tile blocking event arising from $a_{i}^{des}$, it is possible that deterministic growth can occur that results in $E(A_{j\rightarrow i}^{\prime})=E(A_{i})\not\simeq E(A_{j})$ and $E(A_{i\rightarrow j}^{\prime})=E(A_{j})\not\simeq E(A_{i})$. Repulsion-derived tile blocking can thus cause the reasoning in Theorem 3 to fail.
@@
    For any $i\in{1,..,M}$, $G_{S}(S_{i})$ is a tree with at least one branching node, and with branching points in $S_{i}$ labeled $v_{i,n}$. Then, every subtree subshape with origin at some $v_{i,n}$ has a unique shape (Figure 18).
@@
Having expanded our assumptions on our target shape, we wish to consider when repulsion-induced blocking may or may not be significant for our arguments. Consider a sequenced aTAM instance $\mathcal{P}=(A_{0},Q,g)$ that produces a target shape $S$ from starting shape that $S_{0}=E(A_{0})$. The terminal configuration $A_{\zeta}$ of a trajectory of $\mathcal{P}$ has subconfigurations $A_{i}$ such that $E(A_{i})=S_{i}$. Assume that partial sub-configurations $A_{j,t_{s}}$ and $A_{i,t_{s}}=A_{j\rightarrow i,t_{s}}$ could be reached during complete trajectories at some time $t_{s}$. We say that $j$ and $i$ are differentially blocked if, for some single tile configuration $a_{j}$, $A_{j,t_{s}}+a_{j}$ can be reached during a complete trajectory but $A_{i,t_{s}}+a_{j\rightarrow i}$ cannot (or vice versa), due to repulsion-induced blocking (Hence $a_{i}^{des}=a_{j\rightarrow i}$ is the destabilizing tile with coordinate $z_{i}^{des}$). In the following lemma, we show that differential blocking is essential to break the arguments in Theorem 3.
Lemma 4.
@@
Proof.
Consider whether it is possible to detemrinistically produce an assembly with $S_{i}\not\simeq S_{j}$ if $S_{i}$ and $S_{j}$ are anchored to the same glue type. First, if the addition of a tile to partial subconfigurations corresponding to $i$ and $j$ is never blocked due to some repulsive interaction $g<0$, then Theorem 3 is sufficient to argue that $S_{i}\simeq S_{j}$ necessarily, since the assumptions of Definition 7 are a subset of those of Theorem 6 and the contradictions that arise in Theorem 6 remain upheld.
We now allow for non-differential repulsion-induced tile blocking. Consider (without loss of generality) a non-differential repulsion-induced tile block that occurs when trying to add a tile to a partial configuration $A_{i,t_{s}}$, where partial subconfiguration $A_{j,t_{s}}=A_{i\rightarrow j,t_{s}}$ can also be reached in a complete trajectory. This non-differential tile block does not result in partial subconfigurations $A_{j,t_{s}^{\prime}}$ whose equivalents $A_{i,t_{s}^{\prime}}=A_{j\rightarrow i,t_{s}^{\prime}}$ cannot be reached in a complete trajectory. Unless differential tile blocking occurs elsewhere, the assumed properties of the target shape then imply that for all partial subconfigurations of $A_{j}$ that can be reached in a complete trajectory, the equivalent partial subconfiguration of $A_{i}$ can also be reached and the proof of Theorem 6 still applies. Either $S_{i}\simeq S_{j}$ and a violation of assumption 3 of our target shape arises, or growth must be non-deterministic.
Hence, in the absence of differential blocking, each $S_{i}$ must be anchored to a unique glue type, so at least $M$ glue types are required in $A_{0}$. Since each tile type can contain at most $4$ unique glue types, at least $\frac{M}{4}$ tile types, are needed in $A_{0}$ to assemble $S$ deterministically. ∎
Lemma 4 has the following corollary, which states that we would still need $M/4$ tile types if tile blocking only occurs within given subshapes $S_{i}$ rather than in between them.
Corollary 2.
@@
Proof.
We once again assume that $S_{i}$ and $S_{j}$ are anchored to the same glue type. Consider (without loss of generality) a tile blocking event in which $z_{i}^{des}$ only neighbours $S_{i}$, and that the tile is being added to the partial subconfiguration $A_{i,t_{s}}$. Unless a previous differential blocking event has occurred, from the properties of the target shape, the equivalent partial subconfiguration $A_{j,t_{s}}=A_{i\rightarrow j,t_{s}}$ can also be reached in a complete trajectory if assembly is to be oriented-deterministic. $z_{i\rightarrow j}^{des}$ would therefore also experience repulsion-induced blocking, and hence blocking events in which the destabilizing coordinate $z_{i}^{des}$ only neighbours one $S_{i}$ are non-differential. This Corollary then follows directly as a result of Lemma 4.
Lemma 5.
@@
First, assume such a tile blocking event occurs, but $z_{i}^{des}$ is subsequently filled in at some later time (in the trajectory $\Psi^{\prime}$). Then $S_{i}$ would neighbour $S_{k}$, and hence a contradiction arises with either assumption 1 or assumption 2. Hence, we only need to consider differential tile blocks where here $z_{i}^{des}$ is adjacent to $S_{i}$ as well as some $S_{k}$ with $i\neq k$, and $z_{i}^{des}$ remains unoccupied in $A_{\zeta}^{\prime}$, the terminal configuration of $\Psi^{\prime}$.
We now proceed to argue that in this final case, features of $A_{\zeta}^{\prime}$ result in a final shape $S$ that necessarily breaks one of the assumptions 2-6 of Definition 7. $a^{des}_{i}$ neighbours some tile $a^{att}_{j\rightarrow i}$ in $A_{j\rightarrow i,t_{s}}$, onto which it may be attracted, and some other tile $a^{rpl}$ that blocks its addition. Hence, $a^{att}_{j\rightarrow i}$ and $a^{rpl}$ must be of distance $2$ from each other. The assumptions on $S$ that we have laid out forbid points belonging to any two distinct $S_{i}$ and $S_{j}$ from being within distance 2 of each other, with exceptions around the vicinity of junctions where $S_{0}$ meets some $S_{i}$. More specifically, in order to fulfill assumptions 4 and 5, $a^{att}_{j\rightarrow i}$ can only occupy the position $\dot{z}_{i}$ or some neighbour of $\dot{z}_{i}$.
@@
Hence, a contradiction in one of our assumptions arises if a differential block with $z_{i}^{des}$ that neighbours $S_{i}$ and $S_{k}$ with $k\neq i$ occurs. ∎
Theorem 6.
@@
Assumption 7 and determinism imply that $\mathcal{P}$ is oriented-deterministic. As per Lemma 4 and Corollary 2, only differential blocking where $z_{i}^{des}$ neighbours $S_{i}$ and $S_{k}$ with $k\neq i$ will allow $A_{0}$ to require fewer than $\frac{M}{4}$ tile types. However, Lemma 5 shows that such a differential tile block necessarily results in $S$ failing one of assumptions 2- 6 in Definition 7. Hence, a target shape fulfilling the assumptions of Definition 7 cannot be deterministically assembled by $\mathcal{P}$ if $A_{0}$ has fewer than $\frac{M}{4}$ tile types. ∎
@@
The treeangle shape defined in the main text was constructed such that it can be partitioned in a way that satisfies the assumptions of Definition 7. Hence this Theorem can be used in the proof of Lemma 3, taking the place of Theorem 3 which is only valid if $g\geq 0$.
@@
Figure 18: An illustration of assumption 3 in the definition of shapes $S_{i}$ used in Theorem 6. The two shapes are two instances of $S_{i}$, where the lower white tiles connect to some larger $S_{0}$. A single branching point (also the origin tile of the subtree subshapes) is given in orange, while root tiles are outlined in red. The left oriented shape obeys assumption 3, as each subtree subshape (coloured in blue, yellow and green) are distinct, while the right oriented shape violates this assumption as the yellow rooted subtree subshape is equivalent to the green under a $90^{\circ}$ clockwise rotation followed by a $(1,-1)$ translation.
@@
Figure 19: Illustrations of assumptions 4 and 5 in the definition of shapes in Theorem 6. Tiles belonging to $S_{0}$ are in white, and tiles belonging to each $S_{i}$ are assigned a single color. Tiles violating assumptions are outlined in red. a. An illustration of how assumption 4 can be violated. Violation 1 is due to a tile of $S_{j}$ (with $\ddot{z}_{j}=\ddot{z}_{i}$) that is not $\dot{z}_{j}$ being within distance $2$ of $\dot{z}_{i}$, while violation 2 is due to a tile of $S_{k}$ ($\ddot{z}_{k}\neq\ddot{z}_{i}$) being within distance $2$ of an arbitrary tile of $S_{i}$. b. An illustration of how assumption 5 can be violated. Violation 1 is due to a tile of $S_{0}$ being within distance 2 of a tile of $S_{i}$, with both tiles being far from $\ddot{z}_{i}$ and $\dot{z}_{i}$. Violation 2 is due to two tiles (the violating tile and the purple outlined tile) in $S_{0}$ being within distance 2 of $\dot{z}_{i}$. Finally, violation $3$ is due to a tile of $S_{0}$, far from $\ddot{z}_{i}$, being within distance 2 of a neighbour of $\dot{z}_{i}$.
@@
Figure 20: Theorem 6 lays out two possibilities in the case where a tile-blocked coordinate $z^{des}_{i}$ remains unfilled at the end of a complete trajectory. We show here fragments of the terminal configuration $A_{\zeta,c}^{\prime}$. $\mathbf{A_{j\rightarrow i}}$ is given in yellow, $\mathbf{A_{j}}$ is given in blue and destabilizing tiles $\mathbf{z_{i}^{des}}$ are in dotted red,. We illustrate how the two possibilities violate assumptions on $S$ laid out at the beginning of Theorem 6. a. An illustration of how case 1 violates assumptions on $S$. The striped tile can be either a tile of $\mathbf{A_{0}}$ or $\mathbf{\dot{z}_{k}}$. Arrows represent some arbitrary tree-shaped configurations, where arrows of the same color represent configurations with equivalent (up to rotations and translations) shapes. The tile at $z_{j}^{nb}$ cannot be destabilized upon transformation into $z_{i}^{nb}$ as it is too far away from any coordinate of $S_{0}$ or any other $S_{k}$, and hence $z_{i}^{nb}\neq z_{i}^{des}$. Since branches that grow from $z_{i}^{nb}$ are too far away from any $S_{k}$ for $k\neq i$ to be blocked, The branches that grow from $z_{i}^{nb}$ and $z_{j}^{nb}$ must have identical shapes, violating assumptions on the shape $S$. b. An illustration of how case 2 violates assumptions on $S$. Green tiles are $\dot{z}_{k}$ for some $k\neq i,j$. The green arrows represent arbitrary tree-shaped configurations. If $\ddot{z}_{i}$ is branching (Top left), or if $z_{i}^{des}$ neighbours $\dot{z}_{i}$ (Top right), $\{\ddot{z}_{i},\dot{z}_{i}\}$ is a straight line segment of length 2, breaking assumption 6. Otherwise, if the destabilizing tile neighbours $\ddot{z}_{i}$ and a neighbour of $\dot{z}_{i}$ (Bottom left), but its equivalent is present in $S_{j}$, then $S_{j}$ has two coordinates neighbouring $S_{0}$ (Bottom right).